# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v1

@@ III-B Content/Style Modeling
Given an image dataset of two MR contrasts we make the general assumption that there exists an underlying contrast-independent spatial structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images.
Figure 2: A trained content/style model $M$ and the content restoration operator $g_{M}(.;c)$ based on it. (a) T1W/T2W pair and corresponding content maps. Despite general agreement between their content, there is notable discrepancy. (b) Synthetic T2W images generated from 16 style codes grid-sampled from $\mathcal{S}_{2}$, showing a smooth variation in image contrast and a roughly constant anatomical structure. (c) $g_{M}(.;c)$ is applied on 3 cases where $x_{2}^{*}$ was corrupted with $R$=4 undersampling with different center fractions. Sampling more low-frequency lines leads to more accurate style estimate $\hat{s}_{2}$ and thus better image $x_{2}^{\mathrm{restored}}$. Contours in $\mathcal{S}_{2}$ indicate the mean absolute error of the synthetic images from that region.
@@ III-B1 Unpaired pre-training of MUNIT
We formulate our content/style model based on the MUNIT framework [19]. We define “domain” as the set $\mathcal{X}_{i}$ of images of a certain contrast comprising the dataset, where $i\in\{1,2\}$. “Content” $c\in\mathcal{C}$ is defined here as the underlying contrast-independent structure and is represented as a set of spatial maps, whereas “style” $s_{i}\in\mathcal{S}_{i}$ corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g. global effects of acquisition settings, contrast-specific tissue features, etc, and is thus represented as a low-dimensional vector. MUNIT posits the existence of functions $G_{i}^{*}:\mathcal{C}\times\mathcal{S}_{i}\rightarrow\mathcal{X}_{i}$ and their inverses $E_{i}^{*}=(G_{i}^{*})^{-1}$, and learns them jointly via unpaired training given samples from marginal distributions $p(x_{i})$. In practice, the encoder $E_{i}$ is split into content encoder $E_{i}^{c}$ and style encoder $E_{i}^{s}$. Thus, the content/style model is specified as $M=\{E_{1}^{c},E_{2}^{c},E_{1}^{s},E_{2}^{s},G_{1},G_{2}\}$.
The MUNIT loss function comprises of 4 terms:
@@
$\begin{split}\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}=\>&\mathbb{E}_{c_{1}\sim p(c_{1}),s_{2}\sim q(s_{2})}[||c_{1}-E_{1}^{c}(G_{2}(c_{1},s_{2}))||_{1}]+\\ &\mathbb{E}_{c_{2}\sim p(c_{2}),s_{1}\sim q(s_{1})}[||c_{2}-E_{2}^{c}(G_{1}(c_{2},s_{1}))||_{1}],\end{split}$ (9)
@@
$\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}$ is the autoencoder image recovery loss, which promotes preservation of input information in the latent space. $\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}$ and $\mathcal{L}_{\mathrm{style}}^{\mathrm{self}}$ are the content and style recovery losses, respectively, which specialize parts of the latent code into content and style components. $\mathcal{L}_{\mathrm{GAN}}$ is the adversarial loss, which enables unpaired training by enforcing distribution-level similarity between synthetic and real images via discriminators $D_{1}$ and $D_{2}$. $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are hyperparameters.
@@ III-B2 Paired fine-tuning
III-B3 Network architecture and locality bias
@@
Following the original paper [19], our content encoders $E_{i}^{c}$ consist of an input convolutional layer potentially followed by strided downsampling convolutions, and finally a series of residual blocks. Style encoders $E_{i}^{s}$ consist of input and downsampling convolutions followed by adaptive average pooling and a fully-connected layer that outputs the latent vector. Decoders $G_{i}$ follow a similar structure as the content encoders except in reverse. Style is introduced into the decoder via AdaIN operations which modulate the activation maps derived from the content. We note that in this architecture, the content-size-to-image-size ratio reflects the level of spatial structure one expects to be shared between the two domains. This (relative) content size is thus an inductive bias, known as the locality bias [34], built into the model. A model with high locality bias has a large content size, allowing for a richer content that strongly influences the output’s spatial structure, while simultaneously restricting style to learn low-level global features.
III-C Iterative Reconstruction with Content Consistency
@@
Given the content/style model $M$, we consider $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ as reference and target domains, respectively and define the content restoration operator $g_{M}(.;c)$ as
$x^{\mathrm{restored}}_{2}=g_{M}(x^{\mathrm{corrupt}}_{2};c_{1}^{\mathrm{ref}}):=G_{2}(c_{1}^{\mathrm{ref}},E^{s}_{2}(x^{\mathrm{corrupt}}_{2})),$ (14)
@@
where $x^{\mathrm{corrupt}}_{2}$ is the reconstruction corrupted with undersampling artifacts, $x^{\mathrm{restored}}_{2}$ is the restored image, and content $c_{1}^{\mathrm{ref}}=E^{c}_{1}(x^{\mathrm{ref}}_{1})$. This operation restores $x^{\mathrm{corrupt}}_{2}$ by simply replacing its aliased content with the clean content from $x^{\mathrm{ref}}_{1}$, a rule which will later be softened with (15).
@@
Note that this is a radical operation as it discards all structure contained in $x^{\mathrm{corrupt}}_{2}$, retaining only a compact style code $\hat{s}_{2}=E_{2}^{s}(x^{\mathrm{corrupt}}_{2})$. Let $x_{2}^{*}$ be the ground truth reconstruction with content $c_{2}^{*}$ and style $s_{2}^{*}$. The success of the content restoration operation depends on two conditions – (a) $c_{1}^{\mathrm{ref}}$ is close to $c_{2}^{*}$ and (b) $\hat{s}_{2}$ is close to $s_{2}^{*}$. The first condition is roughly satisfied because the model is explicitly trained to minimize content discrepancy, which can be seen in Fig. 2a, but more on this later. On the other hand, it is not obvious that the second condition should hold too and hence deserves a closer look. Assuming a high degree of shared spatial structure between the two domains, the optimal model has high locality bias (based on Section III-B3). Thus, style represents low-level global image features, e.g. contrast variations, as observed in Fig. 2b. It is a well known fact that image contrast is contained prominently in the center of the k-space. Hence, the estimate $\hat{s}_{2}$ can be made arbitrarily close to $s_{2}^{*}$ by sufficiently sampling the k-space center, as shown empirically in Fig. 2c.
@@
3:$x_{2}^{k}\leftarrow A^{H}y$
4:$c^{k}\leftarrow E_{1}^{c}(x^{\mathrm{ref}}_{1})$
@@
Applying data consistency update (3) following $g_{M}(.;c)$ in repetition yields an iterative scheme, where the latter can be viewed as a hard content consistency update. Fig. 2a shows that although $c_{1}^{\mathrm{ref}}$ and $c_{2}^{*}$ are roughly similar, there is notable discrepancy between them. There are at least two possible sources of content discrepancy, which make this phenomenon inevitable – (a) model-related, e.g. practical issues like sub-optimal training or fundamental limits such as irreducible error, and (b) reference image-related, e.g. presence of artifacts independent of the reconstruction. During reconstruction, a $c_{1}^{\mathrm{ref}}$ that conflicts with $c_{2}^{*}$, and with k-space $y$ by extension, would hamper convergence. While model-related discrepancy is partly tackled by the PFT strategy (Section III-B2), we propose a content refinement (CR) step to correct for the remaining discrepancy in the reconstruction stage. The content state is initialized with $c_{1}^{\mathrm{ref}}$ and gradually refined to be consistent with the acquired k-space $y$ as
$c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},E_{2}^{s}(x_{2}^{k}))-y||_{2}^{2},$ (15)
where $\gamma$ is the step size. This update equation mirrors the form of the data consistency update (3), except here the gradient is taken with respect to content through the (non-linear) augmented forward operator $AG_{2}(.)$, which relates the latent content directly to the measured k-space. Hence, by aligning content with k-space explicitly, CR aligns content consistency updates with data consistency updates, correcting for any conflicts between them caused due to content discrepancy. We refer to the resulting algorithm as PnP-MUNIT (Algorithm 1).
IV Experiments
Without loss of generality, we consider the special case of reconstructing T2W scans using T1W references and focus on head applications. In IV-A, we present an analysis of locality bias, convergence, and tuning using BrainWeb-based simulated MR datasets [35]. In IV-B, we benchmark against ablations and closely related methods on NYU fastMRI brain dataset [36, 37] and two in-house multi-coil raw datasets of the brain. Finally, in IV-C, we present a radiological evaluation comparing single- and multi-contrast reconstructions. All software was implemented in PyTorch, and training was performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
IV-A Experiments on Simulated MR Dataset
BrainWeb provides 20 anatomical models of normal brain, each comprised of fuzzy segmentation maps of 12 tissue types. The 20 volumes were first split in 18:1:1 ratio for model training, validation, and reconstruction testing, respectively. T1W/T2W spin-echo scans were simulated using TE/TR values randomly sampled from realistic ranges. For testing, 2D single-coil T2W k-space data was simulated via Fourier transform and 1D Cartesian random sampling.
IV-A1 Interpretation of content size
TABLE I: Reconstruction PSNR (dB) over the BrainWeb test volume (300 slices) at $R$=4. Optimal locality bias is shown in bold. PSNR worse than L1-wavelet CS (16.83$\pm$0.07) is marked as .
Model Config Data Config
RefSize-1 RefSize-2 RefSize-4 RefSize-8
ContentSize-1 27.90$\pm$0.08 21.17$\pm$0.05 19.11$\pm$0.07 16.43$\pm$0.07
ContentSize-2 26.19$\pm$0.08 22.80$\pm$0.08 20.05$\pm$0.08 16.57$\pm$0.08
ContentSize-4 23.98$\pm$0.08 20.88$\pm$0.10 19.44$\pm$0.09 16.88$\pm$0.06
The goal of this experiment was to (a) determine the relationship between the locality bias of the content/style model and the relative spatial resolution of the dataset and (b) explain the effect of reducing the reference image resolution on the reconstruction in terms of locality bias. We simulated 4 datasets, denoted as RefSize-$n$, where $n\in\{1,2,4,8\}$ represents the reference domain downsampling factor. In the $n$=1 case, T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower $1/n$ frequency components. For each dataset, we trained 3 models (for 200k iterations) with locality bias denoted as ContentSize-$m$, where $m\in\{1,2,4\}$ is the content downsampling factor. Table I compares PnP-MUNIT reconstruction quality across these configurations. PFT and CR were disabled here for simplicity. We observe that (a) the optimal locality bias decreases with reference resolution and (b) the reconstruction quality decreases as well, eventually dropping below non-guided reconstruction. Both are explained by the fact that with decrease in reference contrast resolution, the amount of effective content in the dataset decreases, thereby limiting the value of guided reconstruction.
IV-A2 Convergence
We explore here the convergence of PnP-MUNIT by comparing it with two variants. The first uses the content $c_{2}^{*}$ of the ground truth T2W image, thus assuming zero content discrepancy and representing an upper-bound of PnP-MUNIT. The second uses the reference content $c_{1}^{\mathrm{ref}}$ but with CR step disabled, representing a lower-bound of PnP-MUNIT where the non-zero content discrepancy is not corrected. Here as well as in the following simulation experiment, we used the RefSize-1 dataset and the MUNIT-1 model additionally fine-tuned on 2 training volumes for 50k iterations. Fig. 3 shows the convergence curves and intermediate reconstructions. The upper-bound version, given $c_{2}^{*}$, converged in a single iteration, while with $c_{1}^{\mathrm{ref}}$, the ablated version converged equally fast but to a sub-optimal solution. Enabling the CR step closed the gap with the upper-bound, although costing convergence rate.
IV-A3 Tuning the CR step size
The hyperparameter $\gamma$ in Algorithm 1 controls the CR strength. Fig. 4 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal $\gamma$ decreases with the amount and quality of k-space data. This makes sense as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
Figure 5: Evaluation plots over recon-test subset of NYU, LUMC-TRA, and LUMC-COR datasets, consisting of 550, 1366, and 200 slices, respectively.
IV-B Benchmark and Ablation
For benchmarking, we used L1-wavelet CS (CS-WT) and PnP-CNN [20] as our single-contrast baselines. The denoising model of PnP-CNN was defined as a 5-layer network with convolution and ReLU operations and a residual connection after the final layer. We additionally compared against deterministic image translation based on MUNIT defined as $x_{2}^{\mathrm{syn}}:=\mathbb{E}_{s_{2}\sim q(s_{2})}[G_{2}(E_{1}^{c}(x^{\mathrm{ref}}_{1}),s_{2})]$. As guided reconstruction baselines, we used STV-based CS (CS-STV) [6] and image translation combined with L2-regularized reconstruction (Translation + L2 recon) [32] given as $\mathop{\min}_{x_{2}}||Ax_{2}-y||_{2}^{2}+\lambda||x_{2}-x_{2}^{\mathrm{syn}}||_{2}^{2}$. For PnP-MUNIT, we ablated PFT and CR to assess their contribution, and finally, as an upper-bound representing zero content discrepancy, we disabled PFT and CR and used the ideal content $c_{2}^{*}$.

### v2

@@ III-B Content/Style Modeling
Given an image dataset of two MR contrasts we make the general assumption that there exists an underlying contrast-independent structure which, influenced by arbitrary contrast-specific factors, is rendered as the contrast images.
@@ III-B1 Unpaired pre-training of MUNIT
We formulate our content/style model based on the MUNIT framework [22]. We define an image domain as the set $\mathcal{X}_{i}$ of images of a certain contrast comprising the dataset, where $i\in\{1,2\}$. “Content” $c\in\mathcal{C}$ is defined here as the underlying contrast-independent structure and is represented as a set of feature maps, whereas “style” $s_{i}\in\mathcal{S}_{i}$ corresponds to the various modes of variability in one domain which cannot be explained by the other, e.g. global effects of acquisition settings, contrast-specific tissue features, etc., and is represented as a low-dimensional vector. MUNIT posits the existence of functions $G_{i}^{*}:\mathcal{C}\times\mathcal{S}_{i}\rightarrow\mathcal{X}_{i}$ and their inverses $E_{i}^{*}=(G_{i}^{*})^{-1}$, and learns them jointly via unpaired training, given samples from marginal distributions $p(x_{i})$. In practice, the encoder $E_{i}$ is split into content encoder $E_{i}^{c}$ and style encoder $E_{i}^{s}$. Thus, the content/style model is specified as $M=\{E_{1}^{c},E_{2}^{c},E_{1}^{s},E_{2}^{s},G_{1},G_{2}\}$.
The MUNIT loss function is comprised of 4 terms:
@@
$\begin{split}\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}=\>&\mathbb{E}_{c_{1}\sim p(c_{1}),s_{2}\sim q(s_{2})}[||c_{1}-E_{2}^{c}(G_{2}(c_{1},s_{2}))||_{1}]+\\ &\mathbb{E}_{c_{2}\sim p(c_{2}),s_{1}\sim q(s_{1})}[||c_{2}-E_{1}^{c}(G_{1}(c_{2},s_{1}))||_{1}],\end{split}$ (9)
@@
$\mathcal{L}_{\mathrm{image}}^{\mathrm{self}}$ is the image recovery loss, which promotes preservation of the image information in the latent space. $\mathcal{L}_{\mathrm{content}}^{\mathrm{self}}$ and $\mathcal{L}_{\mathrm{style}}^{\mathrm{self}}$ are the content and style recovery losses, respectively, which specialize parts of the latent code into content and style components. $\mathcal{L}_{\mathrm{GAN}}$ is the adversarial loss, which enables unpaired training by enforcing distribution-level similarity between synthetic and real images via discriminators $D_{1}$ and $D_{2}$. $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are hyperparameters.
@@ III-B2 Paired fine-tuning
III-B3 Network architecture and content capacity
@@
Following the original paper [22], our content encoders $E_{i}^{c}$ consist of an input convolutional layer potentially followed by strided downsampling convolutions, and finally a series of residual blocks. Style encoders $E_{i}^{s}$ consist of input and downsampling convolutions followed by adaptive average pooling and a fully-connected layer that outputs the latent vector. Decoders $G_{i}$ follow a similar structure as the content encoders except in reverse. Style is introduced into the decoder via AdaIN operations which modulate the activation maps derived from the content. We observe that in this architecture, the ratio between the content and image resolutions reflects the level of local structure one expects to be shared between the two domains. This (relative) content resolution is thus an inductive bias built into the model, which we refer to as the model’s content capacity. This concept is closely related to the general concept of the locality bias of image-to-image models [43]. A model with high content capacity has a large content resolution, allowing it to learn a rich content representation that strongly influences the output’s structure, which simultaneously restricts the style spaces to learn low-level global features.
Figure 2: A trained content/style model $M$ and the content consistency operator $g_{M}(\cdot;c)$ based on it. Images shown here are from the NYU brain DICOM dataset. Practical details are provided in Section IV. (a) A T1W/T2W image pair and the corresponding content maps. While these two contents generally agree, there is a notable discrepancy. (b) Synthetic T2W images generated from 16 style codes grid-sampled from $\mathcal{S}_{2}$, showing a smooth variation in image contrast and a roughly constant anatomical structure. (c) $g_{M}(\cdot;c)$ is applied on 3 cases where $x_{2}^{*}$ was corrupted with $R$=4 undersampling with different center fractions. Sampling more low-frequency lines leads to more accurate style estimate $\hat{s}_{2}$ and thus a better image $x_{2}^{\mathrm{cc}}$. Contours in $\mathcal{S}_{2}$ indicate the MAE of the synthetic images from that region. The red circle indicate a style estimation NMSE of 0.1.
III-C Iterative Reconstruction using Content Consistency
@@
Given the content/style model $M$, we consider $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ as reference and target domains, respectively, in our guided reconstruction task, and define the content consistency operator $g_{M}(\cdot;c)$ from Eq. (5) as
$x^{\mathrm{cc}}_{2}=g_{M}(x^{\mathrm{us}}_{2};\hat{c}):=G_{2}(\hat{c},E^{s}_{2}(x^{\mathrm{us}}_{2})),$ (14)
@@
where $x^{\mathrm{us}}_{2}$ is the image containing undersampling artifacts, $x^{\mathrm{cc}}_{2}$ is the content-consistent image, and $\hat{c}=E^{c}_{1}(x^{\mathrm{ref}}_{1})$ is the reference content. This operation improves $x^{\mathrm{us}}_{2}$ by simply replacing its aliased content with the content estimated from $x^{\mathrm{ref}}_{1}$, a rule which will later be softened with (16).
@@
Note that this is a radical operation as it discards all structure contained in $x^{\mathrm{us}}_{2}$, retaining only a compact style code $\hat{s}_{2}=E_{2}^{s}(x^{\mathrm{us}}_{2})$. Let $x_{2}^{*}$ be the ground truth reconstruction with content $c^{*}$ and style $s_{2}^{*}$. The success of our content consistency operation depends on two conditions – (a) $\hat{c}$ is close to $c^{*}$ and (b) $\hat{s}_{2}$ is close to $s_{2}^{*}$. The first condition is roughly satisfied, as seen in Fig. 2a, because the model is explicitly trained to minimize content discrepancy, but more on this later. On the other hand, it is not obvious that the second condition should hold too and hence deserves a closer look. Assuming a high degree of shared local structure between the two domains, the optimal model has high content capacity (see Section III-B3). Thus, style would represent mostly low-level global image features, e.g. contrast variations, as indeed observe in Fig. 2b. It is a well known fact that image contrast is contained prominently in the center of the k-space. Hence, the estimate $\hat{s}_{2}$ can be made arbitrarily close to $s_{2}^{*}$ by sufficiently sampling the k-space center, as shown empirically in Fig. 2c.
Applying data consistency update (3) following $g_{M}(\cdot;\hat{c})$ in repetition yields an ISTA-based iterative scheme. Here, our content consistency update complements the data consistency update in the sense that while the latter forces the image estimate to be consistent with the given (measured) k-space data $y$, the former forces it to be consistent with the given (prior) content $\hat{c}$.
@@
3:$x_{2}^{k}\leftarrow A^{H}y$triangleright Initialize Reconstruction
4:$c^{k}\leftarrow E_{1}^{c}(x^{\mathrm{ref}}_{1})$triangleright Initialize Content
@@
The core assumption of our idealized content/style model is that contents $\hat{c}$ and $c^{*}$ are identical. However, in reality, this assumption will not hold, and a discrepancy between the two contents is to be expected. There are two possible sources of this discrepancy – (a) model-related, e.g. fundamental limits such as irreducible error [^1] If perfectly zero error was possible, one could perfectly predict the (distribution of) target image from the reference without any measurement. and practical issues like sub-optimal training, and (b) reference image-related, e.g. presence of artifacts independent of the target reconstruction. This content discrepancy, as observed empirically, is shown in Fig. 2a. During reconstruction, the error in $\hat{c}$ would limit the efficacy of the operator $g_{M}(\cdot;\hat{c})$, affecting reconstruction quality. While model-related discrepancy is partly tackled by PFT (Section III-B2), we now propose a content refinement (CR) process to correct for the remaining discrepancy in the reconstruction stage. Since we have no direct access to the true content $c^{*}$, but only to undersampled k-space measurements $y$, we aim at solving the following minimization problem in CR
$\min_{c}||AG_{2}(c,\hat{s}_{2})-y||_{2}^{2},$ (15)
starting from the initial point $\hat{c}=E^{c}_{1}(x^{\mathrm{ref}}_{1})$ and for a given style estimate $\hat{s}_{2}$. The augmented forward operator $AG_{2}(\cdot)$ is a composition of the linear MRI forward operator $A$ and the non-linear content/style decoder $G_{2}$, and it maps the content domain to the k-space domain. The error between the predicted k-space $AG_{2}(c,\hat{s}_{2})$ and the measured data $y$ serves as a proxy for content discrepancy, which can be computed and minimized during the reconstruction. We approximate the solution with a single gradient descent step
$c^{k}\leftarrow c^{k-1}-\gamma\nabla_{c}||AG_{2}(c^{k-1},\hat{s}_{2}^{k})-y||_{2}^{2},$ (16)
initialized as $c^{0}\leftarrow E^{c}_{1}(x^{\mathrm{ref}}_{1})$ and updated every $k^{\text{th}}$ iteration with step size $\gamma$ following the data consistency and content consistency updates. Hence, by aligning the content $c$ with k-space data $y$ and correcting the discrepancy, the CR module aligns content consistency updates with data consistency updates. With this additional component in place, we obtain our PnP-MUNIT reconstruction algorithm (Algorithm 1). Note that unlike content, the style of the image estimate need not be explicitly aligned with the k-space. The data consistency update applied on the image implicitly “corrects” its style and as the image converges, its style would converge as well.
IV Experimental Setup
IV-A Experiment Design
Without loss of generality, we considered the case of reconstructing T2W scans using T1W references and focused on head applications. We conducted 4 sets of experiments to comprehensively evaluate PnP-MUNIT – (a) simulations, (b) benchmark on NYU DICOM data, (c) benchmark on clinical multi-coil raw data, and (d) radiological evaluation.
TABLE I: Overview of the sequences in the LUMC datasets.
Sequence Params LUMC-TRA LUMC-COR
3D T1W TFE 2D T2W TSE 3D T1W TSE 2D T2W TSE
FA (deg) 8 90 80-90 90
TR (ms) 9.8-9.9 4000-5000 500-800 2000-3500
TE (ms) 4.6 80-100 6.5-16 90-100
ETL 200 14-18 10-13 17-19
Voxel size (mm) 0.98$\times$0.99 $\times$0.91 0.4$\times$0.54 0.59$\times$0.62 $\times$1.19 0.39$\times$0.47
Slice thick. (mm) 3 2
FOV (mm) 238$\times$191 $\times$218 238$\times$190 130$\times$238 $\times$38 130$\times$197
Num slices 50 15